OpenStudy (anonymous):

solve 7 1/2 = x+3 divided by 2

OpenStudy (anonymous):

15=2x+6 ===> x=9/2

OpenStudy (anonymous):

Nope, assuming this is what she meant: 7.5 = (x+3)/2 (15/2) = (x+3)/2 2*(15/2) = x+3 15 = x+3 15-3 = x x=12 plug it in and see 15/2 = 15/2

OpenStudy (anonymous):

or do you mean 7*(1/2) ??

OpenStudy (anonymous):

7.5 = x + 3 divided by 2

OpenStudy (anonymous):

ok then 12 is your x value